### v1

# Mad Libs Are All You Need: Augmenting Cross-Domain Document-Level Event Argument Data

Abstract
Document-Level Event Argument Extraction (DocEAE) is an extremely difficult information extraction problem — with significant limitations in low-resource cross-domain settings. To address this problem, we introduce Mad Lib Aug (MLA), a novel generative DocEAE data augmentation framework. Our approach leverages the intuition that Mad Libs, which are categorically masked documents used as a part of a popular game, can be generated and solved by LLMs to produce data for DocEAE. Using MLA, we achieve a 2.6-point average improvement in overall F1 score. Moreover, this approach achieves a 3.9 and 5.2 point average increase in zero and few-shot event roles compared to augmentation-free baselines across all experiments.
To better facilitate analysis of cross-domain DocEAE, we additionally introduce a new metric, Role-Depth F1 (RDF1), which uses statistical depth to identify roles in the target domain which are semantic outliers with respect to roles observed in the source domain. Our experiments show that MLA augmentation can boost RDF1 performance by an average of 5.85 points compared to non-augmented datasets.
1 Introduction
In recent years, there has been considerable progress in the domain of Event Argument Extraction (EAE), as advances in question answering Du and Cardie (2020), prompt tuning Ma et al. (2022), and semantic graph modeling Yang et al. (2023b) have led to state-of-the-art results on EAE benchmarks. However, some simplifying assumptions made in prior work limit their applicability in a real-world setting, including use of sentence-level data, reliance on large-scale annotation, or that training and testing data come from the same domain and thus share a similar distribution.
Recent works such as Tong et al. (2022); Yang et al. (2023a) have shown that extracting event arguments from long documents given limited training data in a new domain remains extremely challenging. We refer to this problem as few-shot, cross domain (FSCD) document-level EAE (DocEAE) — where document-level extraction requires identifying event roles across multiple sentences.
Figure 1: Mad Libs are a classic game where a humorous story is created by (i) masking phrases from a document (ii) assigning high-level categories to the masked phrases, and (iii) someone who has not read the document fills-in the masked phrases given only the category information. In this paper, we connect the concept of Mad Libs to Event Argument Extraction, showing how generating and solving Mad Libs using LLMs creates high-quality EAE data augmentation.
A key issue in the FSCD DocEAE problem is the need to model 0-shot roles Yang et al. (2023a), i.e. roles which exist in an event’s ontology and thus might appear in the test set, but have no representation in the training data. To illustrate this, consider the DocEE dataset Tong et al. (2022), where each new event in the FSCD training set only has 5 documents available for training. Across all samples representing the Volcano Eruption event, for example, zero documents contain annotation for the role The State of the Volcano (Dormant or Active). This makes this role impossible for token classification models to predict, and requires span extraction models such as question answering to rely on pre-trained knowledge to predict such roles. Few-shot roles in FSCD are also extremely challenging to model and similarly rely on generalizing knowledge learned during pre-training.
One solution to this problem is to generate augmented training samples which increase representation of zero and few-shot roles. However, to the best of our knowledge, no prior work has generated DocEAE training samples in a way which can help resolve the aforementioned issues. For example, many related works in EAE data augmentation such as Gao et al. (2022); Ma et al. (2023); Wang et al. (2023) only focus on sentence-level tasks. The few prior works in DocEAE augmentation do not generate new data samples but rather augment existing samples Liu et al. (2022) or use a pre-trained model to weakly label unannotated corpora Liu et al. (2021).
In this paper, we address this issue by introducing Mad Lib Aug (MLA), a novel, LLM-powered data augmentation framework for document-level EAE. To the best of our knowledge, MLA is the first data augmentation framework for generating data for event-centric document-level argument extraction task. MLA is inspired by the insight that EAE samples can be formulated as a Mad Lib, where the “blanks" found in Mad Libs correspond to roles in an event ontology. We visualize the relationship between Mad Libs and EAE in Figure 1. As shown in Figure 2, we use LLMs to both generate Mad Libs, and then solve them by generating JSON objects, which map event roles to textual arguments. Our method is advantageous in that we (i) formulate EAE data generation in the context of a popular game, Mad Libs, thus leveraging LLMs strong conceptual priors, (ii) our method requires no in-domain data and can generate data for 0-shot roles given only the event title and role name thus facilitating cross-domain EAE, (iii) augmented roles are inserted into Mad Lib “blanks" using string replacement, which is well-suited to extractive tasks. This point is crucial for EAE document generation, as trying to assert specific role arguments occur (or do not occur) during autoregressive generation of a long document can be difficult to perform reliably.
Figure 2: Visualization of our data generation framework. In the Document Generation stage, the LLM treats data generation as a Mad Lib generation task, with instructions to use Mad Lib categories which correspond to event roles. In the role-argument generation stage, the model is instructed to generate a JSON file with the missing information in each (category) slot in the synthesized Mad Lib. An Event Argument Extraction sample can then be produced by replacing the bracketed event role names with their corresponding value in the JSON object.
We evaluate MLA on DocEE, which to the best of our knowledge is the only FSCD EAE dataset for long documents. Our experimental results show that MLA significantly improves performance on zero and few-shot roles while boosting overall F1 score. We additionally show that MLA improves performance on target domain roles which are semantically different from those observed in the source domain. To do so, we introduce a new metric, Role-Depth F1 (RDF1), which uses a statistical depth Seegmiller and Preum (2023) to identify roles in the target domain which are semantic outliers with respect to roles observed in the source domain. For example, the common source-domain role location has semantic similarity with other location-based roles, such as temporary settlement. Conversely, target roles such as maximum wind speed or Magnitude (Tsunami Heights) are domain specific and have little semantic overlap with the source training data. We visualize this phenomena in Figure 3. We argue that RDF1 provides a more useful assessment of cross-domain performance compared to classic F1, which does not capture such semantic nuances.
Our contributions in this study are as follows:
  1. 1.
    We introduce the first document-level EAE data augmentation framework, Mad Lib Aug (MLA). MLA can generate novel documents for DocEAE augmentation and is uniquely suited to facilitate generation of zero-shot, cross-domain event roles.
  2. 2.
    We show that MLA significantly improves performance on a diverse suite of few-shot cross-domain metrics, including a novel EAE evaluation metric, Role-Depth F1 (RDF1). RDF1 characterizes model performance on semantically outlying roles. We show that models trained using data generated with MLA yield significant improvement in RDF1, highlighting the value of MLA in cross-domain DocEAE task.
2 Related Work
2.1 Document-Level Event Argument Extraction
Event Argument Extraction (EAE) is a challenging subtask of the Event Extraction (EE) problem, where the goal is to (i) identify which event(s) occur in the text and (ii) extract event arguments from the text as a structure. Much early work on this problem focuses on sentence level EE, using the popular ACE05  [^1] https://catalog.ldc.upenn.edu/LDC2006T06 dataset. More recently, some work has focused on document-level datasets such as RAMS Ebner et al. (2020), WikiEvents Li et al. (2021), and DocEE Tong et al. (2022). However, RAMS and WikiEvents have significant limitations, as discussed in Tong et al. (2022). RAMS only uses documents with only 5-sentences and WikiEvents contains annotation for only a limited number of documents. DocEE is the first EAE dataset to contain annotation for a large number of full-length documents, with a mean length of $\approx 30$ sentences per document. For this reason, other recent works have chosen to focus their attention on DocEE Yang et al. (2023a). Additionally, DocEE introduces a cross-domain split, which is used to facilitate the cross domain EAE analysis in this work (section 4). Hence, we focus on DocEE in this work.
2.2 Data Augmentation for EAE
To the best of our knowledge, there is no other work that aims to generate synthetic documents for DocEAE. There are however, adjacent works in the realm of EAE data augmentation. In Liu et al. (2021), they use pre-trained EAE models to silver-label (i.e. use a pre-trained model to annotate) unannotated documents as additional data augmentation. In Liu et al. (2022), they use pre-trained language models to augment existing samples by masking annotated argument spans and then generating alternate ones. Gao et al. (2022) do the opposite by first masking unannotated spans, then using a pre-trained T5 model to replace such spans as data augmentation. Wang et al. (2023) perform a reinforcement-learning based solution to sentence-level EE data generation. Contrary to our work, none of the aforementioned methods generate fully novel documents and often rely on augmenting existing samples. Works that augment existing samples have limited control of role-type distribution and are not suited to model 0-shot roles, which is a focal point of our work.
One concurrent work that uses LLMs to generate EAE data for sentence-level tasks is Ma et al. (2023), where they employ LLMs in a multi-step prompt and self-reflect pipeline. However, there are a number of reasons why we cannot apply this work as a baseline. First, their pipeline leverages information that is unavailable for the DocEE dataset, e.g., event definitions, role definitions, and valid entity types per role. Additionally, they provide the LLM with in-domain demonstrations for trigger generation, when our goal is to use no in-domain data for generation. Finally, their proposed pipeline uses significant LLM-based self-reflection. This is a concern as, for document-level tasks (i) self-reflection makes augmentation significantly more expensive when using OpenAI models and (ii) from our experience, self-reflection becomes less effective as document-length increases. Future works may explore how to adapt Ma et al. (2023) to document-level tasks.
2.3 Event Extraction Evaluation
Prior works on EAE in trigger-free datasets use evaluation metrics such as exact match F1 or head noun match F1 Tong et al. (2022). Recently, there have been some advancements in EAE evalulation for trigger-centric datasets. For example, Huang et al. (2023) proposed a new metric (ARG-I+) for trigger-based event extraction to better capture whether a role is associated with the correct trigger. In Li et al. (2022), they explicitly examine performance on long-tail role types. However, to the best of our knowledge, there are no related works on metrics which can measure an EAE model’s performance on semantically outlying roles during cross-domain evaluation. We propose a novel EAE evaluation metric, Role-Depth F1 (RDF1), to address this gap.
3 Methods
In this section we first describe Mad Lib Aug, our LLM-powered data augmentation framework for EAE document generation. Then, we formally introduce our novel metric, Role-Depth F1, which evaluates model performance on semantically outlying role types.
3.1 Mad Lib-based Augmentation
Mad Lib Aug (MLA) leverages the intuition that Mad Libs share a striking resemblance to EAE data. Mad Libs, first released in 1953, are a popular party activity in which a document has various masked phrases, each with a category such as “verb", “car brand", or “type of food". The goal of the game is to choose phrases for each category without knowing the context of the document. The final document is then read aloud, where the chosen category names often produce humorous stories.
The intuition regarding the connection between Mad Libs and EAE data is as follows. If a Mad Lib was written about an event, and used categories which were selected from that event ontology, then a context-aware solution to that Mad Lib would produce EAE annotation. This connection is powerful as LLMs such as GPT-3.5 have significant familiarity with Mad Libs and thus require little to no prompt optimization when requested to generate them. Additionally, Mad Libs are historically narrative, and thus easy to generate using varying styles. In this work, we break EAE data generation down into a two-step framework (i) Mad Lib Generation and (ii) Mad Lib Slot Filling. We describe each step in detail in the following subsections.
3.1.1 Mad Lib Generation
Our Mad Lib generation framework can be found in the top-half of Figure 2. We prompt the LLM to write a Mad Lib about a given <event type> and list each role in the event’s ontology as a category. Thus for each sample, we aim to generate a document which contains every possible role for a given event. Our strategy does not generate documents which include/exclude certain event role subsets. However, we choose this strategy for the following two reasons. (i) In order to perform a fair comparison between augmented and real data in the FSCD setting, we only generate 5 documents per event type. Thus, given our goal is to improve performance on zero and few-shot roles, each document should be as role-dense as possible. (ii) A core challenge of document-level EAE data generation is the issue of generating under-labeled samples. That is, if the desired length of a document about a single event is large, but the number of roles to be included is small, it is very difficult to prevent the LLM from accidentally generating an unintended role. By providing the LLM with all possible target/desired roles as Mad Lib categories, we avoid this issue. To promote high-quality output format, our Mad Lib generation is done using $k=5$ in-context demonstrations, where samples are randomly sampled from a subset of the source domain documents. Specifically, we sample from role-dense documents of manageable length to manage API costs (see Appendix B.2 for details). These randomly sampled documents are then converted into Mad Lib format by replacing argument spans with role names. Note our method has the flexibility to promote document diversity by instructing the model to generate documents in a certain style, such as “newspaper article".
3.1.2 Mad Lib Slot Filling
Once we have generated a Mad Lib, we feed it back into the LLM as shown in the bottom-half of Figure 2. The LLM is instructed to generate a JSON object where the keys are the “categories", or event roles, and the values are the arguments. Note that we choose this format as OpenAI LLMs are able to generate JSON data  [^2] https://platform.openai.com/docs/api-reference/chat/create. We use the $k=5$ documents from the source domain as example data for the LLM in few-shot generation of the JSON object.
We finally do string replacement between the bracketed (category) names in the Mad Lib, and the corresponding key-value pair in the JSON object. The result is an event-centric document with annotation for argument extraction. Note that if during generation the model deviates from generating the exact category name (e.g. generating [the location] instead of [location]), we use SBERT Reimers and Gurevych (2019) to encode the mismatched category name and compute the cosine similarity between it and all possible role names for a given sample. We can reliably match the generated category to the true category using this method and then produce EAE annotation.
More details on Mad Lib Generation and Mad Lib Slot Filling can be found in Appendix B.2. See Appendix C for example augmented data.
Figure 3: Visualization showing how TTE Depth ranks DocEE roles. We find that roles such as “Maximum Wind Speed", “Damaged Crops & Livestock" and “Magnitude (Tsunami Heights)" are outliers compared to roles such as “Influence People" and “Temporary Settlement".
3.2 Role-Depth-F1 (RDF1)
RDF1 aims to evaluate how well cross-domain EAE models are predicting role-types which are semantically different from those observed in the source domain. To compute RDF1, we propose the use of transformer-based text embeddings (TTE) depth Seegmiller and Preum (2023) for identifying the degree to which target domain roles deviate semantically from source domain roles. Formally, given a set of source domain roles, $R_{S}$, and a set of target domain roles, $R_{T}$, TTE depth assigns a score to each target role $r_{t}\in R_{T}$ indicating how well $r_{t}$ represents $R_{S}$.
We embed each role in each set using S-BERT Reimers and Gurevych (2019)333sentence-transformers/all-MiniLM-L6-v2 to get a set of source role embeddings $S$ and a set of target role embeddings $T$. TTE depth scores each target role embedding $t\in T$ according to
$D_{\delta}(t,S):=2-E_{S}[\delta(t,H)]$ (1)
where $H\sim S$ is a random variable with uniform distribution over $S$, and $\delta$ is cosine distance.
TTE depth assigns each target role $r_{t}\in R_{T}$ a score, with higher scores indicating that the target role is representative of the source roles $R_{S}$, and lower scores indicating that the target role is a semantic outlier with respect to the source roles $R_{S}$.
We consider the 25% of target domain roles with the lowest TTE depth scores to be "outlying" roles with respect to the source domain roles. By computing F1 only on these roles, i.e. RDF1, we measure the ability of a FSCD EAE model to perform domain transfer to roles which deviate semantically from the source domain.
In Figure 3, we visualize how depth scores rank all roles in the target domain with respect to the source domain. Notably, we see that Maximum Wind Speed, Damaged Crops & Livestock, and Magnitude (Tsunami heights) are the most semantically outlying, with influence people and temporary settlement being examples of roles closer to source domain roles. This plot justifies the need for RDF1 as a cross-domain evaluation metric, as it follows ones intuition that, for example, temporary settlement may have semantic overlap with source domain concepts like location. Note that this method will inherently filter out evaluation of any overlapping roles between source and target domain.
In DocEE, 13 roles comprise the set of semantic outliers evaluated in RDF1. 12/13 of these roles are few-shot (i.e. < 5 training examples), making RDF1 a more challenging version of the few-shot evaluation subset. We additionally note that the embedding strategy for creating $S$ and $T$ from $R_{S}$ and $R_{T}$ is flexible; event ontologies which include additional information such as role descriptions or valid entity types may be used to enrich the role representations. Since DocEE provides no such information, opting instead for use of semantically-rich role names, the role names themselves are used as $R_{S}$ and $R_{T}$.
4 Experiments
In this section, we describe the details of our experimental setup. Specifically, we introduce the (i) Dataset (ii) Baseline Models, (iii) Baseline Augmentation Methods, and (iv) Evaluation Metrics.
4.1 Dataset
We evaluate our method on the cross-domain split of the DocEE dataset. The DocEE dataset distribution is shown in Table 1. The task is structured to allow pre-training on a large number of source domain events. Then, the goal is to fine-tune this model to a target domain for which there are only very few annotated documents. In this dataset, the target domain is comprised of 10 event types, which are all under the common theme of “Natural Disasters". This grouping of source and target event domains was chosen to reduce overlap between source and target role types Tong et al. (2022). The target and source domains have a small 15% overlap, with common roles such as date, location, and economic loss occurring in both domains. During few-shot cross-domain training, only 5 documents per-event type are provided.
Split # Samples # Events # Argument Instances
Source Train 23,630 49 154,225
Target Train 50 10 308
Target Dev 1,850 10 11,094
Target Test 1,955 10 13,227
Table 1: DocEE data distribution. The source domain contains annotation for over 23k documents across 49 different event types, with over 150k argument instances. The target domain contains annotation for documents with 10 distinct event types, all under the common theme of “Natural Disasters".
4.2 Baseline EAE Models
In this study, we evaluate the impact of EAE data augmentation using the following two EAE models adapted from the DocEE Paper Tong et al. (2022).
  • LongFormer-Seq: This model treats EAE as a token labeling task. Specifically, for a document $D$ with tokens $\{t_{0},\dots,t_{n}\}\in D$, we label argument spans using the label space $\{r_{0},\dots,r_{n}\}\in R$, where a given $r_{i}$ denotes an event role. Given that we are processing long documents, we use the LongFormer-base (149M parameters) Beltagy et al. (2020) architecture, which can process up to 4096 tokens. We use this model as it has been shown to out-perform BERT-based token classifiers on DocEAE Tong et al. (2022).
  • BERT-QA: This model treats EAE as an extractive question-answering task Du and Cardie (2020). In this formulation, the question is the argument role, and the context is the document. The goal is to predict the span(s) $(start_{i},end_{i})$ corresponding to the argument of the role in the document. Note that for this baseline, we use a BERT-base (110M parameters) Devlin et al. (2019) model pre-trained on the SQUAD 2.0 Rajpurkar et al. (2016) question-answering dataset. We choose this baseline as QA-based methods are common in EAE Du and Cardie (2020); Lu et al. (2023) and have been shown to perform well on DocEE in prior work Tong et al. (2022).
We note that the use of two baseline methods is in-line with prior works on data augmentation for Event Extraction Gao et al. (2022). Additionally, we highlight that DocEE is a trigger-free event dataset, meaning that the document is itself event-centric and not grounded to any one trigger mention. This restricts our capacity to evaluate on trigger-dependent EAE baselines Parekh et al. (2023); Huang et al. (2023).
For each experiment, we first fine-tune the model on the source-domain training set, then on the few-shot target dataset. Target domain models are chosen using validation set performance. We train all models on Google Colab using an NVIDIA A100 GPU. Throughout the course of the project we estimate an upper-limit of 75 compute hours. Additional information regarding model hyperparameters, compute resources, and training details can be found in Appendix A.
4.3 Baseline Augmentation Methods
We compare the performance of MLA with a set of suitable augmentation baseline methods. Note that MLA is the first of its kind to generate document-level event arguments. Thus, as baselines we adapt two methods (i) Mask-then-Fill Gao et al. (2022): An augmentation method developed for sentence-level event extraction. This experiment will evaluate how prior models on sentence-level tasks will perform on document-level texts. (ii) Struct2Text: This is a custom baseline we develop which maps event records to documents using LLMs. We base our implementation on a synthetic data generation pipeline provided by LangChain. 444python.langchain.com
  • Mask-then-Fill (MtF): A sentence-level event extraction augmentation strategy which first masks contiguous spans of non-annotated text, then uses T5 Raffel et al. (2020) to fill in the missing spans. Following the implementation of Gao et al. (2022), we first fine-tune T5 on the Gigaword corpus Graff et al. (2003) to familiarize the model with news texts. Next, we split a given training document into event-related and non-event-related fragments, where an event-related fragment is any annotated span. We then mask one non-event-related fragment, and use the fine-tuned T5 model to fill in the masked context.
  • Struct2Text (S2T): We use the data augmentation tool provided by LangChain to (i) generate structured event records in a few-shot setting, then (ii) convert the new event record into a document. The document should contain the exact argument mentioned in the newly generated event record. However, in the case where it fails to do so, we take the closest n-gram in the generated text which matches the argument span in the event record as the annotated span. We use SBERT 555sentence-transformers/all-MiniLM-L6-v2 to encode the spans, and cosine similarity to find the closest matching n-gram. More details of this baseline can be found in Appendix B.3.
For Mad Lib Aug and Struct2Text, we use GPT-3.5 for all text generation. We choose GPT-3.5 due to its capacity to reliably model JSON data, and its relative affordability when compared to GPT-4. All models use default OpenAI parameters with temperature = 1.0. Please refer to Appendix B for additional details on this experiment.
4.4 Evaluation Metrics
We evaluate each model using an exact match F1 score as motivated by prior work Tong et al. (2022). Additionally, we evaluate on a diverse suite of EAE metrics which aim to capture performance on challenging role types. Inspired by Li et al. (2022), for each of the following metrics, we subset samples based on role-type and compute mean F1 score per-role .
  • Role-F1: This metric removes bias towards prediction of common roles. We compute the F1 score of each role type, then report the macro average across all roles. There are 60 total roles in the target domain.
  • 0-Shot F1: This metric identifies performance on roles with no representation in the DocEE target training set. We thus compute F1 score of each role with 0 target domain training samples, then report macro average. There are 9 roles in this subset.
  • Few-Shot F1: This metric identifies performance on roles with low representation in the DocEE target training set. We compute F1 score of each role with less than 5 target domain training samples, then report macro average. There are 31 roles in this subset.
  • RDF1: Our novel metric, RDF1, computes F1 score of each role in the bucket of semantic outliers, then reports the macro average. Statistical depth identifies 13 roles as outliers which are used in this evaluation. See section 3.2 for additional details.
For each metric, we consider the exact match between a set of predicted and ground truth tuples consisting of (document id, role type, argument). We report the mean and standard deviation of three experimental trails with different random seeds for each metric.
5 Results
MLA Boosts Overall Performance:
Our results are shown in Table LABEL:tab:results. We find that for both BERT-QA and LongFormer-Seq, Mad Lib Aug increases overall F1 score by an average of 2.6 F1 points. Interestingly, for BERT-QA we find that training with only synthetic Mad Lib augmentation data, we out-perform the baseline which trains using only real documents across all metrics. This result follows from MLA’s ability to generate high-quality data while maintaining a more even distribution of roles when compared to the raw training data.
We find similar results on the Role-F1 metric, where all models receive a significant boost in performance by employing generative data augmentation. While our results show its advantageous to employ any form of data augmentation to improve Role-F1, MLA stands out as the highest performing model as it increases performance by an average of 4.4 F1 points. Notably, MLA augmentation also has much lower variation across multiple trials when compared to Mask-then-Fill and Struct2Text augmentation on BERT-QA.
MLA Boosts Performance on 0 and Few-Shot Roles:
Our results highlight the significant increase in performance data augmentation provides to 0 and few-shot roles. While both MLA and Struct2Text improve baseline performance on these metrics, MLA is the superior augmentation method on average. When focusing on the Real + Aug MLA experiment, we see that MLA improves baseline performance by an average of 3.9 points and 5.2 points for 0-shot and few-shot roles, respectively. Experiments using Mask-then-Fill augmentation perform on-par or worse than those using no augmentation. This is intuitive as Mask-then-Fill is augmenting existing data, thus it is unable to provide additional information about 0 and few-shot roles.
Data Augmentation Improves Prediction of Semantically Outlying Roles:
Our results highlight that all of our generative data augmentation baselines significantly improve performance on the Role-Depth F1 metric. On BERT-QA, Struct2Text out-performs MLA but with significant variability between experiments. Overall, both methods provide a significant increase in RDF1, which was the most challenging subset for the BERT-QA Real Only baseline. On LongFormer-Seq, MLA out-performs Struct2Text by 0.6 RDF1. Both Struct2Text and MLA provide a 5.4 and 6.0 boost over our no augmentation baseline respectively.
6 Conclusion
In this study, we showcase the capacity of LLMs to generate event-centric document-level argument extraction data. Our method, Mad Lib Aug, produces high-quality DocEAE annotation and significantly improves performance on a variety of cross-domain DocEAE metrics. Mad Lib Aug focuses on generating large categorically templated documents. This method circumvents issues surrounding LLMs failing to follow data generation instructions/constraints, such as including/excluding certain roles in the text. Another key feature of Mad Lib Aug is that it is relatively cost-effective with respect to the number of LLM iterations needed to produce the final output. A future non-template generation approach may require significant self-reflection and/or iteration before producing a document with accurate EAE augmentation. In the future, we hope to explore how Mad Lib Aug can generate documents with greater diversity, from both the Mad Lib Generation perspective (e.g. generating narrative works) and the Mad Lib Slot-Filling perspective (e.g. the Mad Lib solver takes the perspective of a specific person). Additional future work may aim to extend this method to work with trigger-centric event datasets.
7 Limitations
A limitation of our work is use of a closed-source model, GPT-3.5. However, OpenAI models are well-suited to our approach due to their significant capacity to model JSON data  [^6] https://platform.openai.com/docs/api-reference/chat. Future works may explore how to adapt smaller, open-source LLMs for this task as well as the benefit of employing higher-performing models such as GPT-4.
Another limitation of this work is that we did not do an ablation study exploring the following. (i) How varying the number and selection of in-context demonstrations for LLM impacted the overall performance. (ii) How varying the number $n$ of augmented documents per-event type impacted performance. However, we believe that the results in Table LABEL:tab:results clearly demonstrate that MLA produces high-quality DocEAE samples. Future works can explore how to scale this method in effort to further increase the overall performance.
This work is also limited in that we do not explore the capacity of our model to generate large-scale data augmentation with fine-grained control over the role distribution. This was deemed out of scope for this work due to the cost of data generation and the focus on a few-shot task. Additionally, this is an extremely difficult sub-task which requires a more targeted solution. Future works will explore scaling our framework in this way, on both in and cross-domain formatted tasks.
Appendix A Baseline Models
We will make all models available on the Huggingface to ensure reproducibility. Details of our baseline models are found in the following sections.
A.1 BERT-QA
All of the BERT-QA experiments are run using the Huggingface Wolf et al. (2019) Transformers library. We use a BERT-base model (110m parameters) fine-tuned on the SQUAD-v2 dataset 777deepset/bert-base-uncased-squad2 as our base model. Given that BERT has a max token length of 512, we process longer documents in chunks using a stride of 50 tokens. We then fine-tune the model for 3 epochs on the source domain using default Huggingface Trainer parameters  [^8] https://huggingface.co/docs/transformers/main_classes/trainer. During cross-domain fine-tuning, we again train for 3 epochs on target domain data with default parameters and use the model with the best validation loss score for test set prediction. During inference, we take up to 2 predictions per-question, as some documents express more than one argument per role. We use a batch size of 48 for all BERT-QA experiments and train our models on Google Collab use an NVIDIA A100 GPU.
A.2 LongFormer-Seq
All LongFormer-Seq experiments use the LongFormer-base (149M parameters) 999allenai/longformer-base-4096 and are run using the Huggingface Transformers library. As with BERT-QA, we fine-tune the model for 3-epochs on the source training set with default parameters. During cross-domain fine-tuning, we fine-tune for a maximum of 10 epochs using validation F1 score from SeqEval Nakayama (2018) to determine the best model. Note we train this model longer as a new classification head must be learned for cross-domain learning. We use a batch size of 4 for all LongFormer experiments and train our models on Google Collab use an NVIDIA A100 GPU.
A.3 Evaluation Metrics
We use an implementation of F1 score for EAE from the OmniEvent library Peng et al. (2023). When computing F1, since the same argument can be expressed in slightly different ways throughout a document, we normalize the argument string before evaluation using a function provided by the DocEE authors. This makes outputs lowercase, removes articles, white space, and punctuation. For example, If the prediction was (0, date, The 14th of May.), normalization would return (0, date, 14th of may). This is helpful since the same argument can be expressed in many slightly different ways throughout a long document.
Appendix B Data Augmentation Methods
In this section, we discuss various hyperparameters and design choices of our two generative data augmentation models, Mad Lib Aug and Struct2Text.
B.1 Hyperparameters
For each model, we use “gpt-3.5-turbo-0125" from OpenAI. We use temperature=1 with all other parameters left default. We use k=5 in-context learning examples when applicable.
B.2 Mad Lib Aug
B.2.1 Design Choices
Semantic Category Matching:
Mad Lib Aug (MLA) provides a way to generate templates for EAE data, where masked phrases can be string-replaced, removing the problem of reliably generating arguments via LLM instruction. However, given that we are still generating long templates, there is still the issue of reliably generating the masked categories. We find that occasionally, the model will generate a slightly different category, e.g. removing the plurality of the category or adding an additional word. We resolve this by simply computing the cosine similarity (as determined by SBERT 101010sentence-transformers/all-MiniLM-L6-v2) between the all categories and the generated category and returning the best match. We observe that this method reliably returns the correct category.
Few-Shot Sampling Strategy:
When random sampling from the source domain for few-shot Mad Lib generation, we randomly sample from the subset of source domain samples with character length between $\approx 1000-1500$ to manage API costs. From that subset, we focus on sampling from the 20% of documents which are more role-dense.
B.2.2 Prompts
Mad Lib Generation:
You are a creative Mad Lib writer.
Write a madlib for a “[event type]" event in the style of a newspaper article. The number of categories in the text should be exactly [num categories]. A category can occur more than once in the text. Use the following categories for Mad Lib blanks:
Categories: [categories]
Madlib:
Mad Lib Category In-Filling
Fill-in missing information in the document about a "[event type]" event.
For each bracketed [CATEGORY] in the document, fill-in the missing information and add it to a JSON object. Assume all information about the event is already known and publically available.
Make sure the ANSWER written for each [CATEGORY] fits seemlessly into the document. You should output a JSON of general format "CATEGORY: ANSWER".
Document: [document]
JSON:
B.3 Struct2Text
Our Struct2Text baseline follows the synthetic data generation tutorial provided by LangChain 111111python.langchain.com/docs/use_cases/data_generation. This method is based on PyDantic  [^12] https://github.com/pydantic/pydantic and OpenAI Function Calls.
The strategy is to first few-shot generate structured event records. Specifically, we first randomly generate structures as strings from the source domain of the form “role:argument, role:argument, …". We then few-shot prompt the LLM with example structures, while providing it with a PyDantic event schema for our target domain class. The model then generates an argument for each role in the event schema. Finally, the generated event structure is fed to the LangChain OpenAI data generator. Note that we slightly adapt the prompt provided by LangChain in this method to promote generation of documents instead of sentences.
At the sentence level, we observe this method works relatively well with respect to generating texts which follow the event structure (i.e. producing the exact argument found in the event structure). Unfortunately, for long documents, larger schemas, and complex arguments, the LLM tends to deviate away from the structure template. Since we need exact mentions to be found in the text for extractive EAE, we perform semantic n-gram matching with SBERT-encoded n-grams, which identifies the n-gram in the generated document with the highest cosine similarity to the argument in the generated event schema. We empirically choose a threshold of 0.8 similarity for an n-gram to be considered a good match. If the best match is less than 0.8 similarity, we assume the model has not generated any well-matching span and thus do not include annotation for this argument.
We note that on rare occasion, the model can produce structures that fail to parse into the PyDantic framework. In such cases, we simply re-run the model and let the LLM try again. This often quickly resolves itself.
Appendix C Mad Lib Aug Example Data
C.1 Example 1
Generated Mad Lib
[PEOPLE/ORGANIZATION WHO PREDICTED THE DISASTER] successfully forecasted the arrival of [STORM NAME], which hit [STORM HIT LOCATION] on [STORM HIT TIME]. This massive storm, which formed in [STORM FORMATION LOCATION] on [STORM FORMATION TIME], brought torrential rain and wind speeds of up to [MAXIMUM WIND SPEED] mph.
residents were urged to take cover as the storm approached. The storm warning level was raised to [STORM WARNING LEVEL], indicating the severity of the weather system. Many areas experienced [AMOUNT OF PRECIPITATION] of rain, leading to flooding in low-lying areas and road closures.
The storm’s movement speed was recorded at [STORM MOVEMENT SPEED] mph, with a direction that headed towards [STORM DIRECTION]. The storm center passed over [STORM CENTER LOCATION], causing power outages and property damage in its wake.
Despite the efforts of emergency response teams, the storm’s impact left many without power and access to essential services. Residents were advised to stay indoors until the storm passed to ensure their safety. The influence of [STORM NAME] was felt by all in its path, prompting communities to come together in the aftermath to support one another in the recovery process.
Generated JSON Slot Filling
{’PEOPLE/ORGANIZATION WHO
PREDICTED THE DISASTER’: ’The National
Weather Service’, ’STORM NAME’:
’Hurricane Florence’, ’STORM HIT
LOCATION’: ’North Carolina’,
’STORM HIT TIME’: ’September 14, 2018’,
’STORM FORMATION LOCATION’:
’the Atlantic Ocean’,
’STORM FORMATION TIME’:
’September 4, 2018’,
’MAXIMUM WIND SPEED’: ’130’,
’STORM WARNING LEVEL’: ’Category 4’,
’AMOUNT OF PRECIPITATION’:
’over 30 inches’,
’STORM MOVEMENT SPEED’: ’5’,
’STORM DIRECTION’: ’northwest’,
’STORM CENTER LOCATION’: ’Wilmington’}
Final Document
The National Weather Service successfully forecasted the arrival of Hurricane Florence, which hit North Carolina on September 14, 2018. This massive storm, which formed in the Atlantic Ocean on September 4, 2018, brought torrential rain and wind speeds of up to 130 mph.
North Carolina residents were urged to take cover as the storm approached. The storm warning level was raised to Category 4, indicating the severity of the weather system. Many areas experienced over 30 inches of rain, leading to flooding in low-lying areas and road closures.
The storm’s movement speed was recorded at 5 mph, with a direction that headed towards northwest. The storm center passed over Wilmington, causing power outages and property damage in its wake.
Despite the efforts of emergency response teams, the storm’s impact left many without power and access to essential services. Residents were advised to stay indoors until the storm passed to ensure their safety. The influence of Hurricane Florence was felt by all in its path, prompting communities to come together in the aftermath to support one another in the recovery process.
C.2 Example 2
Generated Mad Lib
A devastating earthquake of magnitude [MAGNITUDE] struck the region of [AFFECTED AREA] on [DATE], resulting in catastrophic damage and significant loss of life. The earthquake’s epicenter was located near [EPICENTER], causing widespread destruction across the region.
In the aftermath of the earthquake, [AID AGENCY] quickly mobilized to provide emergency assistance to those affected, delivering [AID SUPPLIES/AMOUNT] in relief supplies to the affected areas. The aid agency also set up temporary settlements to accommodate displaced residents and provide them with basic necessities.
The earthquake led to the destruction of [NUMBER OF DESTROYED BUILDING] buildings, leaving many people homeless. Efforts are underway to rebuild [NUMBER OF REBUILDING HOUSE] homes for those who lost everything in the disaster. Rescue teams worked tirelessly to save [NUMBER OF TRAPPED PEOPLE] people trapped under the debris, managing to rescue [NUMBER OF RESCUED PEOPLE] individuals.
Unfortunately, the earthquake also resulted in a significant number of casualties, with [CASUALTIES AND LOSSES] people losing their lives in the disaster. Many more were injured and had to be evacuated to safety. The economic loss caused by the earthquake is estimated to be around [ECONOMIC LOSS], as the region faces a long road to recovery.
The aftershocks following the earthquake have been frequent, with [NUMBER OF AFTERSHOCKS] tremors further destabilizing the region. The affected area continues to be monitored closely for any potential risks, as residents remain on high alert for any further seismic activity. Our thoughts and prayers are with all those affected by this tragic event.
Generated JSON Slot Filling
{’MAGNITUDE’: ’7.2’,
’AFFECTED AREA’: ’Central Asia’,
’DATE’: ’March 12, 2022’,
’EPICENTER’: ’the town of
Shakhimardan’,
’AID AGENCY’: ’International Red Cross’,
’AID SUPPLIES/AMOUNT’:
’over 100,000 tons’,
’NUMBER OF DESTROYED BUILDING’:
’over 2,000’,
’NUMBER OF REBUILDING HOUSE’: ’1,500’,
’NUMBER OF TRAPPED PEOPLE’: ’50’,
’NUMBER OF RESCUED PEOPLE’: ’38’,
’CASUALTIES AND LOSSES’: ’at least 250’,
’ECONOMIC LOSS’: ’\$50 million’,
’NUMBER OF AFTERSHOCKS’: ’over 100’}
Final Document
A devastating earthquake of magnitude 7.2 struck the region of Central Asia on March 12, 2022, resulting in catastrophic damage and significant loss of life. The earthquake’s epicenter was located near the town of Shakhimardan, causing widespread destruction across the region.
In the aftermath of the earthquake, International Red Cross quickly mobilized to provide emergency assistance to those affected, delivering over 100,000 tons in relief supplies to the affected areas. The aid agency also set up temporary settlements to accommodate displaced residents and provide them with basic necessities.
The earthquake led to the destruction of over 2,000 buildings, leaving many people homeless. Efforts are underway to rebuild 1,500 homes for those who lost everything in the disaster. Rescue teams worked tirelessly to save 50 people trapped under the debris, managing to rescue 38 individuals.
Unfortunately, the earthquake also resulted in a significant number of casualties, with at least 250 people losing their lives in the disaster. Many more were injured and had to be evacuated to safety. The economic loss caused by the earthquake is estimated to be around \$50 million, as the region faces a long road to recovery.
The aftershocks following the earthquake have been frequent, with over 100 tremors further destabilizing the region. The affected area continues to be monitored closely for any potential risks, as residents remain on high alert for any further seismic activity. Our thoughts and prayers are with all those affected by this tragic event.